### v1

@@ Appendix B Lower-density configurations
Figure 10: Normalised confusion matrices illustrating the percentage of predictions for each of the 12 flexion and extension gestures, comparing true labels (y-axis) with predicted labels (x-axis). The matrices show predictions for S1 on the left and S2 on the right, across different data types: a r.m.s. values of simulated bipolar, b r.m.s. values of 8-channel sEMG.
Figure 11: Normalised confusion matrices illustrating the percentage of predictions for each of the 12 flexion and extension gestures, comparing true labels (y-axis) with predicted labels (x-axis). The matrices show predictions for S1 on the left and Subject 2 (S2) on the right, across different data types: a time domain and r.m.s. values of simulated bipolar, b time domain and r.m.s. values of 8-channel sEMG.

### v2

@@ Appendix B Lower-density configurations
Figure 9: Normalised confusion matrices illustrating the percentage of predictions for each of the 12 flexion and extension tasks, comparing true labels (y-axis) with predicted labels (x-axis). The matrices show predictions for participant one on the first row and participant two on the second row, across different data types: RMS values of simulated bipolar (first column), RMS values of 8-channel sEMG (second column).
Figure 10: Normalised confusion matrices illustrating the percentage of predictions for each of the 12 flexion and extension tasks, comparing true labels (y-axis) with predicted labels (x-axis). The matrices show predictions for participant one on the first row and participant two on the second row, across different data types: time domain and RMS values of simulated bipolar (first column), time domain and RMS values of 8-channel sEMG (second column).